# LPR: Large Language Models-Aided Program Reduction

### v2

@@ 5.3.1. Threats to Internal Validity
We mitigate this threat from several perspectives. First, program reduction is a task involving programs distinct from those used in program repair or program synthesis tasks. For instance, LLMs can learn from large datasets about code generation. On the contrary, programs in program reduction tasks are generally large, complex, and most importantly, randomly generated to trigger compiler bugs, with no other specific purpose. Their random and chaotic characteristics make it highly unlikely for LLMs to memorize such disorganized content, thus reducing the risk of data leakage.
Second, we asked LLMs to verify whether they remember any bug-triggering programs from public bug-tracking systems. For instance, we ask LLMs “Do you know the minimal bug-triggering program in GCC’s bugzilla, with id=66186?” and the results confirm that the LLMs do not retain the memory of these programs. Moreover, even in scenarios where the LLMs might coincidentally memorize certain minimal programs, it is improbable for it to link a random-looking, featureless code snippet with a specific memorized program, especially when no explicit bug ID is provided.
Furthermore, Benchmark-JS, one of the benchmark suites used, was created using JIT fuzzing tools by the authors and is not publicly accessible. This exclusivity ensures that the LLMs’ performance on these benchmarks reflects their ability to handle unseen and novel programs, thereby showcasing their effectiveness in managing new challenges without relying on memorized data. This approach significantly mitigates the risk of data leakage and demonstrates LLMs’ capacity for genuine problem-solving and analysis.
@@ 5.3.2. Threats to External Validity
One threat to external validity is the generality of LPR across languages. We evaluate LPR on three prevalent programming languages, namely, C, Rust and JavaScript. We believe these benchmark suites have demonstrated LPR’s high generality on diverse programs. However, without further experiments on a particular language, we still cannot ensure the effectiveness and efficiency of LPR on that language.
An additional threat is the applicability of our approach across different LLMs. To mitigate this threat, we repeat the experiments with other versions of ChatGPT, i.e., gpt-3.5-turbo-1106 and gpt-3.5-turbo-16k-0613. The experimental results show no significant differences from results on gpt-3.5-turbo-0613. Furthermore, as LLMs continue to evolve, we anticipate improvements in both quality and time of code processing.
@@ 6.1. Program Reduction
DDMin (Zeller and Hildebrandt, 2002) initiated the research topic of program reduction. It treats the input as a list of elements, and consistently splits the list into halves. Then it iteratively attempts to reduce the input list by exploring subsets and their complements at varying levels of granularity, transitioning from coarse to fine. Hierarchical Delta Debugging (Misherghi and Su, 2006), short for HDD, parses the program input into a parsing tree, and performs DDMin on each level of the tree structure. Perses (Sun et al., 2018) avoids the generation of syntactically invalid program variants during reduction by formal syntax transformation. Vulcan, further pushes the limit of Perses via identifier/sub-tree replacement and local exhaustive search. All above works are not customized for certain languages, though having high generality, they lack semantic knowledge of a certain language for further reduction.
@@
Distinct from prior work, LPR synergistically combines LLMs and language-generic reduction tool to harness the advantages of both. Language-generic reducers stand out as their remarkable generality across multiple languages, while LLMs excel in further refining the programs with the domain knowledge on certain languages learned from large training sets. Language-specific tools typically demand considerable human effort to design and implement feature-related transformations for reduction, while LPR requires only a few lines of natural languages prompts, significantly reducing the effort involved.
@@ 6.2. LLMs for Software Engineering
Large Language Models (LLMs) have proved their remarkable capability of undertaking multiple text-processing tasks, including source code-related works. Recent works focus on applying LLMs to facilitate software engineering tasks, or assessing the effectiveness, potential and limitations of LLMs on software development and maintenance. Some research (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a) focus on empirically applying LLMs on automatic program repair (APR). Huang et al. (Huang et al., 2023) performed an empirical study on improvement brought by model fine-tuning in APR. Xia et al. (Xia et al., 2023b) thoroughly evaluated 9 state-of-the-art LLMs across multiple datasets and programming languages, and demonstrated that directly applying LLMs has already significantly outperformed all existing APR techniques. Additionally, some works focus on LLMs’ performance w.r.t. code completion, generation and fuzzing (Liu et al., 2023; Zhong and Wang, 2023; Tian et al., 2023a; Deng et al., 2023a, b), by leveraging the code analysis and generation ability of LLMs.
@@ 7. Conclusion
This paper proposes LLMs-aided program reduction (LPR), which is the first approach that leverages LLMs for the program reduction task to the best of our knowledge. By combining the strength of LLMs and existing language-generic program reduction techniques, LPR can perform language-specific transformations to effectively reduce the program while being language-generic (i.e., can be easily applied to a wide range of languages). The evaluation shows that in 50 benchmarks across three programming languages. LPR significantly outperforms Vulcan. Specifically, LPR produces 24.93%, 4.47%, and 11.71% smaller programs on C, Rust, and JavaScript, respectively. Meanwhile, The evaluation also demonstrates that LPR complements Vulcan to some extent. By reducing the outputs of LPR with Vulcan, we attained results that have similar sizes to those of C-Reduce in Benchmark-C. In terms of efficiency, LPR excels in reducing complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks, respectively.

### v3

@@ 5.4.1. Threats to Internal Validity
We mitigate this threat from several perspectives. First, program reduction is a task involving programs distinct from those used in program repair or program synthesis tasks. For instance, LLMs can learn from large datasets about code generation. On the contrary, programs in program reduction tasks are generally large, complex, and most importantly, randomly generated to trigger compiler bugs, with no other specific purpose. Their random and chaotic characteristics make it highly unlikely for LLMs to memorize such disorganized content, thus reducing the risk of data leakage. Moreover, even in scenarios where the LLMs might coincidentally memorize certain minimal programs, it is improbable for it to link a random-looking, featureless code snippet with a specific memorized program, especially when no explicit bug ID is provided. Furthermore, Benchmark-JS, one of the benchmark suites used, was created using JIT fuzzing tools by the authors and is not publicly accessible. This exclusivity ensures that the LLMs’ performance on these benchmarks reflects their ability to handle unseen and novel programs, thereby showcasing their effectiveness in managing new challenges without relying on memorized data. This approach significantly mitigates the risk of data leakage and demonstrates the capacity of LLMs for genuine problem-solving and analysis.
@@ 5.4.2. Threats to External Validity
One threat to external validity is the generalizability of LPR across languages. Although the approach of LPR is language-agnostic, LLMs used by LPR may have limited knowledge of certain languages, which may affect the performance of LPR. To mitigate this threat, we evaluate LPR on three prevalent programming languages, namely, C, Rust and JavaScript. The evaluation results demonstrated the generalizability of LPR on diverse popular programming languages. For languages that are not familiar to LLMs, LPR may still produce reasonable results, since it is shown that LLMs like Codex still perform well in less popular languages like Lua (0.2% in Github) (Cassano et al., 2022). As for a completely new language that LLMs cannot recognize and process, a possible solution is to incorporate the prompts in LPR with few-shot prompting (Ahmed and Devanbu, 2022; Liu et al., 2023b), so that LLMs learn how to recognize and process a new language from the given examples. Further exploration of this approach will be left as future work.
An additional threat is the applicability of our approach across different LLMs. To mitigate this threat, on the same benchmarks used in section 5.2, we repeat the experiments with CodeLlama, another widely-used LLM family. As illustrated in Table 4, the experimental results on two models of CodeLlama, i.e., CodeLlama-13b and CodeLlama-34b, show no significant differences from results on ChatGPT. Furthermore, as LLMs continue to evolve, we anticipate improvements in both quality and time of code processing.
| Perses | Vulcan | LPR with | LPR with CodeLlama | ||
| gpt-3.5-turbo-0613 | 13b | 34b | |||
| Mean | 161.4 | 102.8 | 73.2 ± 3.0 | 73.4 ± 3.5 | 71.3 ± 4.9 |
@@ 6.1. Program Reduction
DDMin (Zeller and Hildebrandt, 2002) initiated the research topic of program reduction. It treats the input as a list of elements, and consistently splits the list into halves. Then it iteratively attempts to reduce the input list by exploring subsets and their complements at varying levels of granularity, transitioning from coarse to fine. Hierarchical Delta Debugging (Misherghi and Su, 2006), short for HDD, parses the program input into a parsing tree, and performs DDMin on each level of the tree structure. Perses (Sun et al., 2018) avoids the generation of syntactically invalid program variants during reduction by formal syntax transformation. Vulcan, further pushes the limit of Perses via identifier/sub-tree replacement and local exhaustive search (Xu et al., 2023). RCC is a compact refreshable caching scheme to speed up program reduction (Tian et al., 2023c). All the above works are not customized for certain languages, though having high generality, they lack semantic knowledge of a certain language for further reduction.
@@
Distinct from prior work, LPR synergistically combines LLMs and language-generic reduction tools to harness the advantages of both. Language-generic reducers stand out for their remarkable generality across multiple languages, while LLMs excel in further refining the programs with the domain knowledge of certain languages learned from large training sets. Language-specific tools typically demand considerable human effort to design and implement feature-related transformations for reduction, while LPR requires only a few lines of natural language prompts, significantly reducing the effort involved.
@@ 6.2. LLMs for Software Engineering
Large Language Models (LLMs) have proved their remarkable capability of undertaking multiple text-processing tasks, including source code-related works. Recent works focus on applying LLMs to facilitate software engineering tasks, or assessing the effectiveness, potential and limitations of LLMs on software development and maintenance. Some research (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a) focus on empirically applying LLMs on automatic program repair (APR). Huang et al. (Huang et al., 2023) performed an empirical study on improvement brought by model fine-tuning in APR. Xia et al. (Xia et al., 2023b) thoroughly evaluated 9 state-of-the-art LLMs across multiple datasets and programming languages, and demonstrated that directly applying LLMs has already significantly outperformed all existing APR techniques. Additionally, some works focus on LLMs’ performance w.r.t. code completion, generation and fuzzing (Liu et al., 2023a; Zhong and Wang, 2023; Tian et al., 2023a; Deng et al., 2023), by leveraging the code analysis and generation ability of LLMs.
@@ 7. Conclusion
This paper proposes LLMs-aided program reduction (LPR), which is the first approach that leverages LLMs for the program reduction task to the best of our knowledge. By combining the strength of LLMs and existing language-generic program reduction techniques, LPR can perform language-specific transformations to effectively reduce the program while being language-generic (i.e., can be easily applied to a wide range of languages). The evaluation shows that in 50 benchmarks across three programming languages. LPR significantly outperforms Vulcan. Specifically, LPR produces 24.93%, 4.47%, and 11.71% smaller programs on C, Rust, and JavaScript, respectively. Meanwhile, The evaluation also demonstrates that LPR complements Vulcan to some extent. By reducing the outputs of LPR with Vulcan, we attained results that have similar sizes to those of C-Reduce in Benchmark-C. In terms of efficiency perceived by users, LPR excels in reducing complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks, respectively.
8. Data Availability
For replication, we have implemented LPR and released it publicly (Zhang et al., 2024).
Acknowledgements.
We thank all the anonymous reviewers in ISSTA’24 for their insightful feedback and comments. This research is partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant, a project under WHJIL, and CFI-JELF Project #40736.References
- (1)
- ope (2023a) 2023a. OpenAI API. Retrieved 2023-11-20 from https://platform.openai.com/docs/overview
- ope (2023b) 2023b. OpenAI API: N. Retrieved 2023-11-20 from https://platform.openai.com/docs/api-reference/chat/create#chat-create-n
- ope (2023c) 2023c. OpenAI API: Temperature. Retrieved 2023-11-20 from https://platform.openai.com/docs/api-reference/chat/create#chat-create-temperature
- Ahmed and Devanbu (2022) Toufique Ahmed and Premkumar Devanbu. 2022. Few-shot training llms for project-specific code-summarization. In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering. 1–5. https://doi.org/10.48550/arXiv.2207.04237
- Cassano et al. (2022) Federico Cassano, John Gouwar, Daniel Nguyen, Sydney Nguyen, Luna Phipps-Costin, Donald Pinckney, Ming-Ho Yee, Yangtian Zi, Carolyn Jane Anderson, Molly Q Feldman, et al. 2022. Multipl-e: A scalable and extensible approach to benchmarking neural code generation. arXiv preprint arXiv:2208.08227 (2022). https://doi.org/10.48550/arXiv.2208.08227
- Deng et al. (2023) Yinlin Deng, Chunqiu Steven Xia, Haoran Peng, Chenyuan Yang, and Lingming Zhang. 2023. Large language models are zero-shot fuzzers: Fuzzing deep-learning libraries via large language models. In Proceedings of the 32nd ACM SIGSOFT international symposium on software testing and analysis. 423–435. https://doi.org/10.1145/3597926.3598067
- Donaldson and MacIver (2021) Alastair Donaldson and David MacIver. 2021. Test Case Reduction: Beyond Bugs. Retrieved May 29, 2023 from https://blog.sigplan.org/2021/05/25/test-case-reduction-beyond-bugs
- Donaldson et al. (2021) Alastair F Donaldson, Paul Thomson, Vasyl Teliman, Stefano Milizia, André Perez Maselco, and Antoni Karpiński. 2021. Test-case reduction and deduplication almost for free with transformation-based compiler testing. In Proceedings of the 42nd ACM SIGPLAN International Conference on Programming Language Design and Implementation. 1017–1032. https://doi.org/10.1145/3453483.3454092
- Fan et al. (2023b) A. Fan, B. Gokkaya, M. Harman, M. Lyubarskiy, S. Sengupta, S. Yoo, and J. M. Zhang. 2023b. Large Language Models for Software Engineering: Survey and Open Problems. In 2023 IEEE/ACM International Conference on Software Engineering: Future of Software Engineering (ICSE-FoSE). IEEE Computer Society, Los Alamitos, CA, USA, 31–53. https://doi.org/10.1109/ICSE-FoSE59343.2023.00008
- Fan et al. (2023a) Zhiyu Fan, Xiang Gao, Martin Mirchev, Abhik Roychoudhury, and Shin Hwei Tan. 2023a. Automated repair of programs from large language models. In 2023 IEEE/ACM 45th International Conference on Software Engineering (ICSE). IEEE, 1469–1481. https://doi.org/10.1109/ICSE48619.2023.00128
- Gu (2023) Qiuhan Gu. 2023. LLM-Based Code Generation Method for Golang Compiler Testing. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra, Kelly Blincoe, and Paolo Tonella (Eds.). ACM, 2201–2203. https://doi.org/10.1145/3611643.3617850
- Heo et al. (2018) Kihong Heo, Woosuk Lee, Pardis Pashakhanloo, and Mayur Naik. 2018. Effective Program Debloating via Reinforcement Learning. In Proceedings of the 2018 ACM SIGSAC Conference on Computer and Communications Security (Toronto, Canada) (CCS ’18). Association for Computing Machinery, New York, NY, USA, 380–394. https://doi.org/10.1145/3243734.3243838
- Huang et al. (2023) Kai Huang, Xiangxin Meng, Jian Zhang, Yang Liu, Wenjie Wang, Shuhao Li, and Yuqing Zhang. 2023. An Empirical Study on Fine-Tuning Large Language Models of Code for Automated Program Repair. In 2023 38th IEEE/ACM International Conference on Automated Software Engineering (ASE). IEEE Computer Society, 1162–1174. https://doi.org/10.1109/ASE56229.2023.00181
- Jain et al. (2022) Naman Jain, Skanda Vaidyanath, Arun Shankar Iyer, Nagarajan Natarajan, Suresh Parthasarathy, Sriram K. Rajamani, and Rahul Sharma. 2022. Jigsaw: Large Language Models meet Program Synthesis. In 44th IEEE/ACM 44th International Conference on Software Engineering, ICSE 2022, Pittsburgh, PA, USA, May 25-27, 2022. ACM, 1219–1231. https://doi.org/10.1145/3510003.3510203
- Kalhauge and Palsberg (2019) Christian Gram Kalhauge and Jens Palsberg. 2019. Binary reduction of dependency graphs. In Proceedings of the 2019 27th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering. 556–566. https://doi.org/10.1145/3338906.3338956
- Kalhauge and Palsberg (2021) Christian Gram Kalhauge and Jens Palsberg. 2021. Logical bytecode reduction. In Proceedings of the 42nd ACM SIGPLAN International Conference on Programming Language Design and Implementation. 1003–1016. https://doi.org/10.1145/3453483.3454091
- Le et al. (2014) Vu Le, Mehrdad Afshari, and Zhendong Su. 2014. Compiler validation via equivalence modulo inputs. ACM Sigplan Notices 49, 6 (2014), 216–226. https://doi.org/10.1145/2594291.2594334
- Lecoeur et al. (2023) Bastien Lecoeur, Hasan Mohsin, and Alastair F. Donaldson. 2023. Program Reconditioning: Avoiding Undefined Behaviour When Finding and Reducing Compiler Bugs. Proc. ACM Program. Lang. 7, PLDI, Article 180 (jun 2023), 25 pages. https://doi.org/10.1145/3591294
- Li et al. (2023a) Junyi Li, Xiaoxue Cheng, Xin Zhao, Jian-Yun Nie, and Ji-Rong Wen. 2023a. HaluEval: A Large-Scale Hallucination Evaluation Benchmark for Large Language Models. In The 2023 Conference on Empirical Methods in Natural Language Processing. https://doi.org/10.48550/arXiv.2305.11747
- Li et al. (2023b) Tsz On Li, Wenxi Zong, Yibo Wang, Haoye Tian, Ying Wang, Shing-Chi Cheung, and Jeff Kramer. 2023b. Nuances are the Key: Unlocking ChatGPT to Find Failure-Inducing Tests with Differential Prompting. In 38th IEEE/ACM International Conference on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 14–26. https://doi.org/10.1109/ASE56229.2023.00089
- Liu et al. (2023a) Jiawei Liu, Chunqiu Steven Xia, Yuyao Wang, and Lingming Zhang. 2023a. Is your code generated by chatgpt really correct? rigorous evaluation of large language models for code generation. arXiv preprint arXiv:2305.01210 (2023). https://doi.org/10.48550/arXiv.2305.01210
- Liu et al. (2023b) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2023b. Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing. Comput. Surveys 55, 9 (2023), 1–35. https://doi.org/10.1145/3560815
- Livinskii et al. (2020) Vsevolod Livinskii, Dmitry Babokin, and John Regehr. 2020. Random testing for C and C++ compilers with YARPGen. Proceedings of the ACM on Programming Languages 4, OOPSLA (2020), 1–25. https://doi.org/10.1145/1993498.1993532
- LLVM (2000) LLVM. 2000. LibTooling. https://clang.llvm.org/docs/LibTooling.html Accessed: 2023-04-30.
- Misherghi and Su (2006) Ghassan Misherghi and Zhendong Su. 2006. HDD: hierarchical delta debugging. In Proceedings of the 28th International Conference on Software Engineering. 142–151. https://doi.org/10.1145/1134285.1134307
- Niemetz and Biere (2013) Aina Niemetz and Armin Biere. 2013. ddSMT: a delta debugger for the SMT-LIB v2 format. In Proceedings of the 11th International Workshop on Satisfiability Modulo Theories, SMT. 8–9.
- Regehr et al. (2012) John Regehr, Yang Chen, Pascal Cuoq, Eric Eide, Chucky Ellison, and Xuejun Yang. 2012. Test-case reduction for C compiler bugs. In Proceedings of the 33rd ACM SIGPLAN Conference on Programming Language Design and Implementation. 335–346. https://doi.org/10.1145/2254064.2254104
- Regehr (2012) John et al. Regehr. 2012. C-Reduce. Retrieved 2023-11-26 from https://github.com/csmith-project/creduce
- Shaham et al. (2022) Uri Shaham, Elad Segal, Maor Ivgi, Avia Efrat, Ori Yoran, Adi Haviv, Ankit Gupta, Wenhan Xiong, Mor Geva, Jonathan Berant, and Omer Levy. 2022. SCROLLS: Standardized CompaRison Over Long Language Sequences. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, EMNLP 2022, Abu Dhabi, United Arab Emirates, December 7-11, 2022, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, 12007–12021. https://doi.org/10.18653/V1/2022.EMNLP-MAIN.823
- Shi et al. (2023) Freda Shi, Xinyun Chen, Kanishka Misra, Nathan Scales, David Dohan, Ed H Chi, Nathanael Schärli, and Denny Zhou. 2023. Large language models can be easily distracted by irrelevant context. In International Conference on Machine Learning. PMLR, 31210–31227. https://doi.org/10.5555/3618408.3619699
- Sun et al. (2016a) Chengnian Sun, Vu Le, and Zhendong Su. 2016a. Finding compiler bugs via live code mutation. In Proceedings of the 2016 ACM SIGPLAN International Conference on Object-Oriented Programming, Systems, Languages, and Applications. 849–863. https://doi.org/10.1145/2983990.2984038
- Sun et al. (2016b) Chengnian Sun, Vu Le, Qirun Zhang, and Zhendong Su. 2016b. Toward understanding compiler bugs in GCC and LLVM. In Proceedings of the 25th International Symposium on Software Testing and Analysis. 294–305. https://doi.org/10.1145/2931037.2931074
- Sun et al. (2018) Chengnian Sun, Yuanbo Li, Qirun Zhang, Tianxiao Gu, and Zhendong Su. 2018. Perses: Syntax-guided program reduction. In Proceedings of the 40th International Conference on Software Engineering. 361–371. https://doi.org/10.1145/3180155.3180236
- Sun et al. (2023) Maolin Sun, Yibiao Yang, Yang Wang, Ming Wen, Haoxiang Jia, and Yuming Zhou. 2023. SMT Solver Validation Empowered by Large Pre-Trained Language Models. In 38th IEEE/ACM International Conference on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 1288–1300. https://doi.org/10.1109/ASE56229.2023.00180
- Tian et al. (2023a) Haoye Tian, Weiqi Lu, Tsz On Li, Xunzhu Tang, Shing-Chi Cheung, Jacques Klein, and Tegawendé F Bissyandé. 2023a. Is ChatGPT the Ultimate Programming Assistant–How far is it? arXiv preprint arXiv:2304.11938 (2023). https://doi.org/10.48550/arXiv.2304.11938
- Tian et al. (2023d) Jia Le Tian, Mengxiao Zhang, Zhenyang Xu, Yongqiang Tian, Yiwen Dong, and Chengnian Sun. 2023d. Ad Hoc Syntax-Guided Program Reduction. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra, Kelly Blincoe, and Paolo Tonella (Eds.). ACM, 2137–2141. https://doi.org/10.1145/3611643.3613101
- Tian et al. (2023b) Yongqiang Tian, Zhenyang Xu, Yiwen Dong, Chengnian Sun, and Shing-Chi Cheung. 2023b. Revisiting the Evaluation of Deep Learning-Based Compiler Testing. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence, IJCAI 2023, 19th-25th August 2023, Macao, SAR, China. ijcai.org, 4873–4882. https://doi.org/10.24963/IJCAI.2023/542
- Tian et al. (2023c) Yongqiang Tian, Xueyan Zhang, Yiwen Dong, Zhenyang Xu, Mengxiao Zhang, Yu Jiang, Shing-Chi Cheung, and Chengnian Sun. 2023c. On the Caching Schemes to Speed Up Program Reduction. ACM Trans. Softw. Eng. Methodol. 33, 1, Article 17 (nov 2023), 30 pages. https://doi.org/10.1145/3617172
- Wang et al. (2021) Guancheng Wang, Ruobing Shen, Junjie Chen, Yingfei Xiong, and Lu Zhang. 2021. Probabilistic Delta debugging. In Proceedings of the 29th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering. 881–892. https://doi.org/10.1145/3468264.3468625
- Wang et al. (2024) Junjie Wang, Yuchao Huang, Chunyang Chen, Zhe Liu, Song Wang, and Qing Wang. 2024. Software Testing with Large Language Models: Survey, Landscape, and Vision. https://doi.org/10.48550/arXiv.2307.07221 arXiv:2307.07221 [cs.SE]
- Wang et al. (2023b) Junjie Wang, Zhiyi Zhang, Shuang Liu, Xiaoning Du, and Junjie Chen. 2023b. FuzzJIT: Oracle-Enhanced Fuzzing for JavaScript Engine JIT Compiler. In USENIX Security Symposium. USENIX.
- Wang et al. (2023a) Theodore Luo Wang, Yongqiang Tian, Yiwen Dong, Zhenyang Xu, and Chengnian Sun. 2023a. Compilation Consistency Modulo Debug Information. In Proceedings of the 28th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (Vancouver, BC, Canada) (ASPLOS 2023). Association for Computing Machinery, New York, NY, USA, 146–158. https://doi.org/10.1145/3575693.3575740
- Wei et al. (2023) Yuxiang Wei, Chunqiu Steven Xia, and Lingming Zhang. 2023. Copiloting the Copilots: Fusing Large Language Models with Completion Engines for Automated Program Repair. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra, Kelly Blincoe, and Paolo Tonella (Eds.). ACM, 172–184. https://doi.org/10.1145/3611643.3616271
- Wu et al. (2023) Yi Wu, Nan Jiang, Hung Viet Pham, Thibaud Lutellier, Jordan Davis, Lin Tan, Petr Babkin, and Sameena Shah. 2023. How Effective Are Neural Networks for Fixing Security Vulnerabilities. In Proceedings of the 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis (¡conf-loc¿, ¡city¿Seattle¡/city¿, ¡state¿WA¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (ISSTA 2023). Association for Computing Machinery, New York, NY, USA, 1282–1294. https://doi.org/10.1145/3597926.3598135
- Xia et al. (2023a) Chunqiu Steven Xia, Yifeng Ding, and Lingming Zhang. 2023a. Revisiting the Plastic Surgery Hypothesis via Large Language Models. arXiv preprint arXiv:2303.10494 (2023). https://doi.org/10.48550/ARXIV.2303.10494
- Xia et al. (2023b) Chunqiu Steven Xia, Yuxiang Wei, and Lingming Zhang. 2023b. Automated program repair in the era of large pre-trained language models. In Proceedings of the 45th International Conference on Software Engineering (ICSE 2023). Association for Computing Machinery. https://doi.org/10.1109/ICSE48619.2023.00129
- Xia and Zhang (2023) Chunqiu Steven Xia and Lingming Zhang. 2023. Keep the Conversation Going: Fixing 162 out of 337 bugs for $0.42 each using ChatGPT. arXiv preprint arXiv:2304.00385 (2023). https://doi.org/10.48550/arXiv.2304.00385
- Xu et al. (2023) Zhenyang Xu, Yongqiang Tian, Mengxiao Zhang, Gaosen Zhao, Yu Jiang, and Chengnian Sun. 2023. Pushing the Limit of 1-Minimality of Language-Agnostic Program Reduction. Proceedings of the ACM on Programming Languages 7, OOPSLA1 (2023), 636–664. https://doi.org/10.1145/3586049
- Yang et al. (2011) Xuejun Yang, Yang Chen, Eric Eide, and John Regehr. 2011. Finding and understanding bugs in C compilers. In Proceedings of the 32nd ACM SIGPLAN Conference on Programming Language Design and Implementation. 283–294. https://doi.org/10.1145/1993316.1993532
- Zeller and Hildebrandt (2002) Andreas Zeller and Ralf Hildebrandt. 2002. Simplifying and isolating failure-inducing input. IEEE Transactions on Software Engineering 28, 2 (2002), 183–200. https://doi.org/10.1109/32.988498
- Zhang et al. (2024) Mengxiao Zhang, Yongqiang Tian, Zhenyang Xu, Yiwen Dong, Shin Hwei Tan, and Chengnian Sun. 2024. Artifact for ”LPR: Large Language Models-Aided Program Reduction”. https://doi.org/10.5281/zenodo.10901714
- Zhang et al. (2023) Mengxiao Zhang, Zhenyang Xu, Yongqiang Tian, Yu Jiang, and Chengnian Sun. 2023. PPR: Pairwise Program Reduction. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering. 338–349.
- Zhong and Wang (2023) Li Zhong and Zilong Wang. 2023. A study on robustness and reliability of large language model code generation. arXiv preprint arXiv:2308.10335 (2023). https://doi.org/10.48550/arXiv.2308.10335